The perimeter of a rectangle is 36 inches, and the area is 77 square inches. Find the length and width of the rectangle
Answer by mananth(16946)

Perimeter = 2(L+W)
2(L+W) = 36
L+W= 18
Area = 77 in^2
area = L*W
L*W= 77
L= 77/W
substitute L in first equation
77/W + W = 18
multiply by W
77+W^2= 18W
W^2-18W+77=0
W^2-11W-7W+77=0
W(W-11)-7(W-11) =0
(W-11)(W-7)=0
therefore the dimensions are 7& 11 m